# Vizing’s Theorem in Near-Linear Time

### v1

@@ 1.2 Technical Overview
Collecting U-Fans in General Graphs: We now discuss the generalization of our scheme above to non-bipartite graphs. Existence of odd cycles in non-bipartite graphs implies that we can no longer assign a color type $\{c_{u},c_{v}\}$ to an uncolored edge $(u,v)$ for $c_{u}\in\mathsf{miss}_{\chi}(u),c_{v}\in\mathsf{miss}_{\chi}(v)$, and hope that flipping the $\{c_{u},c_{v}\}$-alternating path from $u$ allows us to color the edge $(u,v)$ with $c_{v}$ (because the path may end at $v$, and thus after flipping it $c_{v}$ will no longer be missing from $v$). This is where Vizing fans and Vizing chains come into play: in non-bipartite graphs, a color type of an uncolored edge $(u,v)$ is $\{c_{u},\gamma_{u,v}\}$ where $c_{u}$ is an arbitrary color in $\in\mathsf{miss}_{\chi}(u)$ but $\gamma_{u,v}$ is determined by the Vizing fan around $u$ and the alternating path that we take for coloring $(u,v)$ (or the Vizing chain of $(u,v)$). Thus, while as before we can switch $c_{u}$ with some fixed color $\alpha$, it is unclear how to flip alternating paths to change $\gamma_{u,v}$ to $\beta$ also, in order to popularize the edge $(u,v)$ to be of some designated type $\{\alpha,\beta\}$ —without damaging another popularized edge as a result.
@@
Let us make this a bit more precise. Fix an arbitrary color $\alpha$ and consider the set $U_{\alpha}$ of all the uncolored edges $(u,v)$ such that $\alpha\in\mathsf{miss}_{\chi}(u)$ and the respective Vizing chain is of type $\{\alpha,\cdot\}$. Also, assume there are $m_{\alpha}$ edges colored $\alpha$ under $\chi$. If most $\{\alpha,\cdot\}$-Vizing chains have length larger than $\Omega(m_{\alpha}/|U_{\alpha}|)$, then on average, two Vizing chains should meet within the first $O(m_{\alpha}/|U_{\alpha}|)$ steps; in this case, we can repeatedly pick two intersecting Vizing chains and create a u-fan by shifting their initial uncolored edges to the intersection of these Vizing chains. Given the length of the chains, this takes $O(m_{\alpha}/|U_{\alpha}|)$ time. Otherwise, the average cost of applying Vizing’s color extension procedure is $O(m_{\alpha}/|U_{\alpha}|)$, and in this case we can directly color all those edges in $O(m_{\alpha})$ time. Summing over all different $\alpha\in[\Delta+1]$ gives a near-linear runtime. The above argument will be detailed in the proof of Lemma 4.1 in Section 5.
The above discussion leaves out various technical challenges. For instance, moving around uncolored edges as described above breaks the assumption that the uncolored edges form a matching. Handling this requires dedicating different colors from $\mathsf{miss}_{\chi}(u)$ for every uncolored edge incident on a vertex $u$. This is formalized via the notion of separability in Section 3.1. Additionally, we have ignored all algorithmic aspects of (efficiently) finding pairs of intersecting Vizing chains, as well as the corner cases of Vizing fan intersections and fan-chain intersections. We defer the discussions on these details to the main body of the paper.
2 Preliminaries
We begin by describing the basic notation used throughout the paper in Section 2.1 and then give the classic definitions of Vizing fans and Vizing chains in Section 2.2. In Section 3, we introduce the new definitions that provide the basic building blocks for our algorithm.
2.1 Notation
Let $G=(V,E)$ be graph on $n$ vertices with $m$ edges and maximum degree $\Delta$ and let $\chi:E\longrightarrow[\Delta+1]\cup\{\bot\}$ be a (partial) $(\Delta+1)$-edge coloring of $G$. We refer to edges $e\in E$ with $\chi(e)=\bot$ as uncolored. Given a vertex $u\in V$, we denote the set of colors that are not assigned to any edge incident on $u$ by $\mathsf{miss}_{\chi}(u)$. We sometimes refer to $\mathsf{miss}_{\chi}(u)$ as the palette of $u$. We say that the colors in $\mathsf{miss}_{\chi}(u)$ are missing (or available) at $u$.

### v3

@@ 1.2 Technical Overview
Collecting U-Fans in General Graphs: We now discuss the generalization of our scheme above to non-bipartite graphs. Existence of odd cycles in non-bipartite graphs implies that we can no longer assign a color type $\{c_{u},c_{v}\}$ to an uncolored edge $(u,v)$ for $c_{u}\in\mathsf{miss}_{\chi}(u),c_{v}\in\mathsf{miss}_{\chi}(v)$, and hope that flipping the $\{c_{u},c_{v}\}$-alternating path from $u$ allows us to color the edge $(u,v)$ with $c_{v}$ (because the path may end at $v$, and thus after flipping it $c_{v}$ will no longer be missing from $v$). This is where Vizing fans and Vizing chains come into play: in non-bipartite graphs, a color type of an uncolored edge $(u,v)$ is $\{c_{u},\gamma_{u,v}\}$ where $c_{u}$ is an arbitrary color in $\mathsf{miss}_{\chi}(u)$ but $\gamma_{u,v}$ is determined by the Vizing fan around $u$ and the alternating path that we take for coloring $(u,v)$ (or the Vizing chain of $(u,v)$). Thus, while as before we can switch $c_{u}$ with some fixed color $\alpha$, it is unclear how to flip alternating paths to change $\gamma_{u,v}$ to $\beta$ also, in order to popularize the edge $(u,v)$ to be of some designated type $\{\alpha,\beta\}$, without damaging another popularized edge as a result.
@@
[…]
Figure 3: In this picture, we have two different uncolored edges $(u_{1},v_{1}),(u_{2},v_{2})$ such that $\alpha\in\mathsf{miss}_{\chi}(u_{1})\cap\mathsf{miss}_{\chi}(u_{2})$, and their Vizing chains first intersect at edge $(x,y)$ which currently has color $\alpha$ under $\chi$. Then we can rotate their Vizing fans and flip part of their Vizing chains to shift $(u_{1},v_{1}),(u_{2},v_{2})$ to $(w_{1},x),(w_{2},x)$ respectively to form a u-fan; note that $\alpha\in\mathsf{miss}_{\chi}(w_{1})\cap\mathsf{miss}_{\chi}(w_{2})$ after this shifting procedure.
Let us make this a bit more precise. Fix an arbitrary color $\alpha$ and consider the set $U_{\alpha}$ of all the uncolored edges $(u,v)$ such that $\alpha\in\mathsf{miss}_{\chi}(u)$ and the respective Vizing chain is of type $\{\alpha,\cdot\}$. Also, assume there are $m_{\alpha}$ edges colored $\alpha$ under $\chi$. If most $\{\alpha,\cdot\}$-Vizing chains have length larger than $\Omega(m_{\alpha}/|U_{\alpha}|)$, then on average, two Vizing chains should meet within the first $O(m_{\alpha}/|U_{\alpha}|)$ steps; in this case, we can repeatedly pick two intersecting Vizing chains and create a u-fan by shifting their initial uncolored edges to the intersection of these Vizing chains; see Figure 3 for illustration. Given the length of the chains, this takes $O(m_{\alpha}/|U_{\alpha}|)$ time. Otherwise, the average cost of applying Vizing’s color extension procedure is $O(m_{\alpha}/|U_{\alpha}|)$, and in this case we can directly color all those edges in $O(m_{\alpha})$ time. Summing over all different $\alpha\in[\Delta+1]$ gives a near-linear runtime. The above argument will be detailed in the proof of Lemma 6.2 in Section 7.
The above discussion leaves out various technical challenges. For instance, moving around uncolored edges as described above breaks the assumption that the uncolored edges form a matching. Handling this requires dedicating different colors from $\mathsf{miss}_{\chi}(u)$ for every uncolored edge incident on a vertex $u$. This is formalized via the notion of separability in Section 5.1. Additionally, we have ignored all algorithmic aspects of (efficiently) finding pairs of intersecting Vizing chains, as well as the corner cases of Vizing fan intersections and fan-chain intersections. We defer the discussions on these details to the actual proofs in subsequent sections.
2 Basic Notation
Let $G=(V,E)$ be graph on $n$ vertices with $m$ edges and maximum degree $\Delta$ and let $\chi:E\rightarrow[\Delta+1]\cup\{\bot\}$ be a (partial) $(\Delta+1)$-edge coloring of $G$. We refer to edges $e\in E$ with $\chi(e)=\bot$ as uncolored. Given a vertex $u\in V$, we denote the set of colors that are not assigned to any edge incident on $u$ by $\mathsf{miss}_{\chi}(u)$. We sometimes refer to $\mathsf{miss}_{\chi}(u)$ as the palette of $u$. We say that the colors in $\mathsf{miss}_{\chi}(u)$ are missing (or available) at $u$.
@@ 2 Basic Notation
Consider a set $U\subseteq E$ of edges that are uncolored under $\chi$, i.e., $\chi(e)=\bot$ for all $e\in U$. We use the phrase “extending $\chi$ to $U$ to mean the following: Modify $\chi$ so as to ensure that $\chi(e)\neq\bot$ for all $e\in U$, without creating any new uncolored edges. When the set $U$ consists of a single edge $e$ (i.e., when $U=\{e\}$), we use the phrase “extending $\chi$ to the edge $e$ instead of “extending $\chi$ to $U$.